OpenStudy (anonymous):

Find the sum twelve and three-eighths plus two and one-fifth. Be sure to reduce to simplest form. fourteen and four-thirteenths fourteen and twenty-three fortieths twelve and twenty-three fortieths fourteen and four-fifths

OpenStudy (driftracer305):

is it \[12\frac{ 3 }{ 8 } + 2\frac{ 1 }{ 5 }\]??

OpenStudy (kva1992):

115/8 = 14 3/8

OpenStudy (kva1992):

and yes drift thats how i solved it

OpenStudy (driftracer305):

actually @kva1992 .......... u are very wrong

OpenStudy (anonymous):

http://gyazo.com/211310b0842ad68451e2841ad974f7e1

OpenStudy (driftracer305):

if thats the quaestion...... then further solving it equals to 583/40

OpenStudy (kva1992):

but you would have to reduce

OpenStudy (kva1992):

wait snaps did i add wrong?

OpenStudy (driftracer305):

yes....... and it reduces to \[14\frac{ 23 }{ 40 }\]......

OpenStudy (kva1992):

opps yeah i did i forgot to add 1

OpenStudy (kva1992):

thanks for catching that

OpenStudy (kva1992):

my bad Mikayla

OpenStudy (driftracer305):

sure......... no prb...... @MikaylaSkyy .... hope u get it

OpenStudy (kva1992):

you can only do soo much in your head

OpenStudy (anonymous):

Yea thanks guys c:

OpenStudy (driftracer305):

thats y i took some time.........did all mentally....lol

OpenStudy (kva1992):

haha well glad one of us got it right lol

OpenStudy (driftracer305):

well i had to............

OpenStudy (kva1992):

haha no option to get it wrong lol